OpenStudy (maddiie):

Find the solution to each system by the substitution method. 2x-5y=-1 x-3y=-3

OpenStudy (amistre64):

2x-5y=-1 x-3y=-3 (-2) 2x -5y =-1 -2x +6y = 6 ------------ y = 5

OpenStudy (anonymous):

smells like teen spirit here

OpenStudy (amistre64):

x = 12; y=5

OpenStudy (amistre64):

to sub tho is to find a 'value' from one equation and substitute it in the other to get a solid value

OpenStudy (amistre64):

its very labor intensive .... but works

OpenStudy (anonymous):

or use a matrix

OpenStudy (anonymous):

and matlab

OpenStudy (immanuelv):

cannot use matrix on this

OpenStudy (anonymous):

I know, in regard to it being labor intensive

OpenStudy (amistre64):

2x-5y=-1 x-3y=-3 <-- x = 3y-3 2x-5y=-1 2(3y-3)-5y=-1 6y -6-5y =-1 y = 5 .... lol

OpenStudy (amistre64):

x = 3y-3 x = 3(5)-3 x = -15+3 x = -12

OpenStudy (amistre64):

very much labor inducing lol

OpenStudy (anonymous):

haha for real

OpenStudy (anonymous):

ive seen enough facoring type questions.. i want to see integration

OpenStudy (maddiie):

This was very confusing.

OpenStudy (amistre64):

its like using a chain rule for solving equations....

OpenStudy (amistre64):

x = 'value'; use x in other equation: y = 'real value'; use in other equation x = real value

OpenStudy (maddiie):

Another on one. 10x+3y=-8 2x+y=2

OpenStudy (maddiie):

I'm going to try to do it on my end

OpenStudy (anonymous):

good girl

OpenStudy (anonymous):

a lot of ppl here don't try

OpenStudy (amistre64):

y = -2x+2 .... use it in other equaiton 10x +3(-2x+2) = -8 10x -6x +6 = -8 16x = -14 x = -7/8

OpenStudy (amistre64):

use x value in other equation to solve for y y = -2(-7/8)+2 y = 7/4 +2 y = 15/4

OpenStudy (maddiie):

Thanks. But you have to try because if not you aren't learning anything just getting answers.

OpenStudy (amistre64):

10x+3y=-8 2x+y=2 (-3) 10x+3y=-8 -6x -3y=-6 ----------- 4x = -14 x = -7/2 ..... lol

OpenStudy (anonymous):

actually, we will be doing the elimination

OpenStudy (amistre64):

i added wrong last time lol 10-6 = 16? what a farce

OpenStudy (amistre64):

2x -7/2 = 2 4x -7 = 4 4x = 11 x = 11/4

OpenStudy (amistre64):

just shoot me lol

OpenStudy (amistre64):

2(-7/2) +y = 2 -7 +y = 2 y = 9 .... and that my final offer lol

OpenStudy (amistre64):

thatll teach ya to ask 2 questions in one post ;)

OpenStudy (maddiie):

lol... I'm trying to figure it out and you guys keep going. AHHHHH

OpenStudy (anonymous):

2x-5y=-1 (x-3y=-3)*-2 ----> -2x+6y=6 then add, y=5 2x-5y=-1 2x-5(5)=-1 2x-25=-1 2x=25-1 2x=24 x=12

OpenStudy (gw2011):

The answer is x=12 and y=5

OpenStudy (anonymous):

i hope this helps.

OpenStudy (amistre64):

dbl check 10x+3y=-8 10(-7/2) +3(9) = -8 -35 + 27 = -8 -8 = -8 ...... thats good 2x+y=2 2(-7/2) +(9) = 2 -7 +9 = 2 2 = 2 ....... thats good x = -7/2 and y = 9

OpenStudy (amistre64):

yeah i had a -12 up top eh.. silly me lol

OpenStudy (anonymous):

is this your answers?